# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

### v2

@@ 3. Reference Subject Extraction
[…]
@@ 5. Compute revised attention:
Appendix B Baselines
In both qualitative and quantitative comparisons, we only compared our FreeGraftor with training-free approaches and publicly available zero-shot models. Detailed descriptions of these methods are provided below.
  • IP-Adapter(Ye et al., 2023) introduces a lightweight module that projects visual features from an image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
  • MS-Diffusion(Wang et al., 2024a) tackles multi-subject personalization by using layout guidance and grounding tokens to ensure each subject’s details are preserved and positioned in defined regions. Its specialized cross-attention mechanism effectively minimizes subject conflicts while achieving high visual fidelity in a zero-shot setting.
  • OmniGen(Xiao et al., 2024) unifies diverse image generation tasks, including text-to-image synthesis, editing, subject-driven generation and even classical vision operations—within a single diffusion framework based solely on a VAE and transformer. Trained on a large-scale unified dataset, it avoids additional encoders and transfers knowledge robustly across tasks and domains.
@@ Appendix B Baselines
  • FreeCustom(Ding et al., 2024) provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism that integrates features from different reference images. With its weighted masking strategy to suppress extraneous details, it efficiently composes high-quality images without further training.
@@
  • DiptychPrompting(Shin et al., 2024) reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the attention between panels, it captures fine-grained subject details in a zero-shot manner.
Appendix C Additional Experiments
C.1. Efficiency Analysis
Table 3. Efficiency Comparison of Subject-Driven Generation Methods
@@ C.1. Efficiency Analysis
Method Base Model Time (s) Memory (MiB)
FreeCustom Stable Diffusion v1.5 22.02 14670
IP-Adapter FLUX.1-dev 9.58 40788
MS-Diffusion SDXL $\mathbf{8.89}$ 14654
OmniGen Phi-3 46.13 $\mathbf{12638}$
DiptychPrompting FLUX.1-dev 52.11 46858
Ours FLUX.1-dev 42.62 41106
We compare the efficiency of FreeGraftor with test-time subject-driven generation methods (including zero-shot and training-free approaches) on an NVIDIA A40 GPU. All methods generate 512×512 resolution images under identical conditions, with measurements of peak GPU memory consumption and inference time.
As shown in Table 3, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms DiptychPrompting in both time and memory efficiency, while demonstrating comparable memory usage to IP-Adapter. Across all evaluated methods, OmniGen achieves the lowest memory footprint but requires substantially longer inference time. MS-Diffusion attains the fastest inference speed and second-lowest memory consumption after OmniGen, though these efficiency gains come at the expense of reduced subject fidelity and degraded visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
Method Text-alignment Image-alignment
FreeCustom 1.813 1.638
IP-Adapter 3.090 2.183
MS-Diffusion 3.458 2.905
OmniGen 3.905 3.368
DiptychPrompting 4.633 4.280
Ours $\mathbf{4.815}$ $\mathbf{4.773}$
Table 4. Results of user study. Our method is the most preferred by users, achieving the highest ratings in both text and image alignment.
C.2. User Study
We conduct a user study assessing human preference of text-to-image correspondence and subject identity consistency in generated results. Participants are presented with images synthesized by competing methods alongside their corresponding textual prompts and reference subjects. For text alignment evaluation, participants rate the semantic consistency between generated images and prompts on a 5-point Likert scale (1: minimal correspondence, 5: perfect alignment). For image alignment, they evaluate the visual similarity between generated subjects and reference counterparts using an identical scale, with higher scores indicating stronger identity preservation. The study involves a comprehensive evaluation comprising 40 participants, each assessing 60 randomly sampled images. As detailed in Table 4, our method demonstrated significant superiority over baselines in both metrics, validating its effectiveness in balancing textual faithfulness and subject fidelity.
C.3. More visualization results
Figure 10 presents additional qualitative results comparing our method with the baselines. As shown in these results, our method generates text-aligned novel scenes while maintaining subject fidelity. Our FreeGraftor consistently preserves pixel-level visual details, even for challenging text and patterns, highlighting its effectiveness in subject-driven generation.
@@ C.3. More visualization results
Extended single-subject generation results with various prompts are visualized in Figure 11. Our method demonstrates comprehensive prompt adaptability, dynamically modulating subject attributes, poses, and scene configurations to synthesize high-fidelity images that achieve both visual harmony and precise semantic alignment with input textual descriptions.
Figure 10. More qualitative comparisons
Figure 11. Single-subject generation results
Appendix D Limitations
While FreeGraftor achieves robust subject-driven generation in most scenarios, its performance partially depends on the reliability of external models (e.g., GroundingDINO, SAM, LaMa) during the collage construction stage. For instance, rare failures in subject detection, imperfect segmentation masks, or residual artifacts from inpainting could propagate to subsequent stages, occasionally leading to suboptimal initialization. Additionally, the current implementation builds upon the FLUX.1 base model, which demands significant GPU memory (e.g., 40+ GB for 1024×1024 resolution), limiting accessibility for users with constrained hardware. While this choice ensures compatibility with state-of-the-art architectures, adapting our framework to lighter MM-DiT variants or exploring model distillation and quantization techniques could broaden its practical applicability.
Appendix E Social Impacts
The proposed FreeGraftor framework has the potential to democratize personalized content creation by enabling users to synthesize high-fidelity images of specific subjects without requiring extensive training. This capability could empower creators in fields such as digital art, advertising, and education to rapidly prototype visual concepts while preserving intricate details (e.g., logos, textures), thereby lowering barriers to professional-grade content generation. However, the ease of transferring visual identities also raises concerns about misuse, such as generating deceptive imagery for disinformation campaigns or unauthorized replication of copyrighted subjects. Additionally, while our method avoids the computational costs of per-subject fine-tuning, the reliance on large pre-trained diffusion models still necessitates significant energy consumption during inference, highlighting the need for ongoing research into efficient deployment strategies. We advocate for ethical guidelines that balance creative freedom with safeguards against malicious applications, such as watermarking synthesized content or restricting access to sensitive reference data.

### v3

@@ 3. Reference Subject Extraction
$B^{\text{ref}}\leftarrow\text{GroundingDINO}(I^{\text{ref}},\text{"subject"})$
@@ 6. $I^{\text{gen}}\leftarrow\text{SAFG}(\epsilon^{\text{init}},\mathcal{F}^{\text{ref}},P)$ ;
Figure 4. Illustration of the Semantic-aware Feature Grafting (SAFG) Module. For each patch in the reference image, it first establishes semantic correspondence in the generated image via feature matching. The module then (1) concatenates the key and value pairs of the reference patch with those of the corresponding generated patch, while (2) copying and applying the position embedding from the generated patch to the reference patch’s key. This mechanism enables effective positional information sharing between corresponding patches while maintaining semantic alignment.
@@ 5. Compute revised attention:
3.2. Semantic-Aware Feature Grafting
Inspired by appearance transfer techniques(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024), we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Fig. 4).
3.2.1. Semantic Matching
3.2.2. Position-Constrained Attention Fusion
After obtaining the filtered reference mask $M^{\text{ref}}$, we transfer features from the retained reference patches. A naive approach is to directly replace the features of the corresponding patches. However, this may introduce artifacts along object boundaries or even cause image distortion (as shown in Fig. 7). Unlike U-Net, which models positional relationships through convolutional operations, MM-DiT distinguishes patch proximity and orientation via position embeddings (PE). This enables us to align reference patches with their target positions in the generated image by editing position embeddings.
(9) $K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
For each retained reference patch $i$, we extract the corresponding position embedding $PE_{m(i)}^{\text{img, gen}}$ based on the matching relationship $m(i)$. The new position embedding is then formed by concatenating the following embeddings:
The revised attention computation is formulated as:
Here, $A^{+}$ represents the updated attention scores, and $H_{\text{out}}^{+}$ denotes the output of the attention layer. By integrating semantic matching and the sharing of position embeddings, our SAFG module effectively binds features between matching patches in the reference and generated images, facilitating cross-image feature grafting.
3.3. Structure-Consistent Initialization
Attention-based feature migration replicates the appearance of reference subjects (e.g., color, texture) but may not preserve structural details (e.g., shape, body proportions). Moreover, Gaussian noise initialization introduces significant randomness, leading to discrepancies in shape and size between generated subjects and those in the reference image, potentially causing semantic matching failures. To maintain structural integrity and enhance the robustness of semantic matching, we propose a structure-consistent initialization strategy. This involves creating a collage containing the reference subject and then employing an inversion technique to derive the initial noise required for the generation phase.
As illustrated in Fig. 3, given a text prompt (e.g., ”a cat sitting on a chair wearing a birthday hat”), we first use a base text-to-image model to generate a template image that faithfully represents the described scene. To preserve the structure of the reference subject, we replace the corresponding subject in the template image with the reference subject. Specifically, a grounding model localizes the target subject in the template, a segmentation model extracts its mask, and an inpainting model removes the subject from the template. The reference subject is then cropped (using the same grounding and segmentation models), resized, and pasted into the erased region to form a collage. This collage replaces the original reference image as the new reference. Finally, we invert this collage to obtain the initial noise for the generation phase and record the diffusion trajectory during inversion to facilitate feature grafting.
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps(Balaji et al., 2022; Alaluf et al., 2024), we reduce the injection of reference features during the early steps while retaining them in the later steps to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
(14) ${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage. This avoids simultaneous processing of multiple reference images and eliminates additional computational overhead.
4. Experiments
4.1. Experimental Setup
4.1.1. Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO(Liu et al., 2024) to localize subjects, SAM(Kirillov et al., 2023) to segment subjects, and LaMa(Suvorov et al., 2022) to erase the original subject in the template image. FireFlow(Deng et al., 2024) is employed for rectified flow inversion. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$. All experiments are conducted on an NVIDIA L20 GPU.
4.1.2. Baselines
We compare our method with zero-shot methods (BLIP-Diffusion, IP-Adapter(Ye et al., 2023), MS-Diffusion(Wang et al., 2024a), UNO(Wu et al., 2025), DreamO(Mou et al., 2025)) and training-free methods (FreeCustom(Ding et al., 2024), DiptychPrompting(Shin et al., 2024), LatentUnfold(Kang et al., 2025)). Detailed descriptions of these methods are provided below.
@@ 4.1.2. Baselines
  • FreeCustom(Ding et al., 2024) provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism along with a weighted mask strategy so that the generated image attends to each reference concept. With its weighted masking strategy to suppress extraneous details. This design efficiently preserves the semantic content of each input concept within context without further training.
  • BLIP-Diffusion(Li et al., 2023) trains a multimodal encoder to produce a unified text-aligned visual representation of the target subject. During training, it learns to leverage this subject representation so the diffusion model can generate new renditions of the subject from text prompts.
  • MS-Diffusion(Wang et al., 2024a) injects reference images and explicit grounding tokens using a CLIP-based feature resampler to preserve fine details for each reference subject. Using the provided layout as guidance, it applies a multi-subject cross-attention mechanism so that each subject’s conditioning affects a specific region of the image, which harmonizes the composition of all subjects while respecting the text prompt.
  • IP-Adapter(Ye et al., 2023) introduces a lightweight module that projects visual features from a CLIP image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
  • UNO(Wu et al., 2025) first uses the inherent in-context generation ability of MM-DiTs to synthesize large-scale, high-consistency multi-subject paired data. Then it trains a “subject-to-image” model (conditioned on multiple images) with progressive cross-modal alignment and a universal rotary positional embedding. This pipeline yields a model that maintains high consistency while supporting control in both single- and multi-subject generation tasks.
  • DreamO(Mou et al., 2025) is a unified image customization framework built on a MM-DiT. DreamO is trained on a large mixed dataset covering diverse tasks (identity, subject, style, background, etc.), and uses techniques to integrate multiple conditions. It employs a feature-routing constraint that queries the relevant information from each reference, and a placeholder strategy that assigns each condition a dedicated position in the input. DreamO can perform many customization tasks with high quality and flexibly combine different types of control conditions in one model.
@@
  • DiptychPrompting(Shin et al., 2024) reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the cross-attention between the two halves, DiptychPrompting precisely aligns the generated subject with the reference.
  • LatentUnfold(Kang et al., 2025) reframes subject generation as grid-based image completion: the reference images are simply tiled in a mosaic layout to “activate” the model’s latent identity features. This simple “free lunch” approach is enhanced with a cascade attention design and a meta-prompting technique to boost fidelity.
4.1.3. Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench(Ruiz et al., 2023), Mix-of-Show(Gu et al., 2023) and CustomConcept101(Kumari et al., 2023) for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench(Ruiz et al., 2023), with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP(Radford et al., 2021) and ImageReward(Xu et al., 2024) metrics. For image alignment, we use GroundedSAM(Ren et al., 2024) to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP(Radford et al., 2021) and DINOv2(Oquab et al., 2023).
4.2. Quantitative Results
As shown in Table 1, methods using smaller base models (FreeCustom and BLIP-Diffusion) exhibit limitations in both image and text alignments. MS-Diffusion and IP-Adapter achieve lower image alignment than UNO and DreamO, suggesting that reliance solely on CLIP-based external encoders is insufficient for preserving intricate subject details. Training-free methods (DiptychPrompting and LatentUnfold) underperform in image alignment compared to trained zero-shot alternatives (UNO and DreamO), which indicates that the inherent context learning capacity of the base model itself becomes a limiting factor. Crucially, our method achieves the highest alignment across all metrics, demonstrating superior subject fidelity and textual adherence, validating its effectiveness for subject-driven generation.
Image-alignment Text-alignment Efficiency
@@ 4.2. Quantitative Results
Method Base Model CLIP-I DINO CLIP-T ImageReward Time (s) Memory (MiB)
FreeCustom SD v1.5 0.8308 0.6107 0.3246 0.6223 14.93 12068.73
BLIP-Diffusion SD v1.5 0.8621 0.7364 0.2949 0.2360 1.44 3641.84
MS-Diffusion SDXL 0.9023 0.7977 0.3254 1.3405 4.02 12268.41
IP-Adapter FLUX.1-dev 0.8920 0.7696 0.3048 0.7444 7.45 38690.51
UNO FLUX.1-dev 0.9069 0.8237 0.3151 1.2788 14.02 35248.43
DreamO FLUX.1-dev 0.9159 0.8400 0.3217 1.6234 12.15 34431.61
DiptychPrompting FLUX.1-dev 0.8924 0.7971 0.3291 1.5728 43.27 44928.55
LatentUnfold FLUX.1-dev 0.8741 0.7755 0.3170 1.3929 104.90 35402.22
Ours FLUX.1-dev $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$ 40.93 28993.33
In additon, we compare the efficiency of FreeGraftor with test-time subject-driven generation methods (including zero-shot and training-free approaches) on an NVIDIA L20 GPU. All methods generate 512×512 resolution images under identical conditions, with measurements of peak GPU memory consumption and inference time. As shown in Table 1, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms all other approaches in terms of memory usage, while also achieving shorter inference time compared to the other two training-free methods (DiptychPrompting and LatentUnfold). Compared to zero-shot methods, our approach is also competitive. While IP-Adapter, UNO, and DreamO have faster inference speeds than FreeGraftor, they consume more GPU memory. Due to their smaller parameter sizes, FreeCustom, BLIP Diffusion, and MS-Diffusion achieve high efficiency in both time and memory, but this comes at the cost of reduced subject consistency and lower visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
Figure 6. Generated results with different text prompts under the same reference image.
4.3. Qualitative Results
4.3.1. Single-Subject Generation
Fig. 5 presents the results for single-subject generation using FreeGraftor and other methods. FreeCustom, BLIP-Diffusion and IP-Adapter fail to retain fine visual details of the reference subject effectively. While other baselines partially recover reference characteristics, they struggle with fine-grained fidelity. FreeGraftor achieves pixel-level detail preservation, such as logos on the can, numbers on the clock, and text on the teddy bear. For non-rigid subjects (e.g., dogs, plush toys), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
@@ 4.3.1. Single-Subject Generation
Extended single-subject generation results with various prompts are visualized in Fig. 6. Our method demonstrates comprehensive prompt adaptability, dynamically modulating subject attributes, poses, and scene configurations to synthesize high-fidelity images that achieve both visual harmony and precise semantic alignment with input textual descriptions.
4.3.2. Multi-Subject Generation
As shown in Fig. 7, FreeCustom and MS-Diffusion struggle to achieve sufficient subject consistency. While UNO and DreamO demonstrate higher fidelity, they suffer from subject omission (e.g., the barn and the table) and attribute confusion (e.g., confusion between the cat and the dog). FreeGraftor preserves all reference details while enabling flexible text-guided control, highlighting its effectiveness in multi-subject generation.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
w/o Structure-consistent Initialization 0.8529 0.7138 0.3236 1.6320
w/o Feature Grafting 0.8463 0.6572 0.3306 1.6434
w/o Semantic Matching 0.9504 0.8974 0.3232 1.5258
w/ Feature Replacement 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
4.4. Ablation Study
4.4.1. Component Ablation
To validate the effectiveness of each component in our method, we compare several ablation variants, as shown in Fig. 8. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors and patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, there is no visual correlation between the generated and reference subjects. (c) If semantic matching is omitted and reference keys/values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table 2 confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as both are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
Figure 9. Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.:
Figure 10. Generation results under different dropout intensity ($\omega$) settings. Higher $\omega$ improves text alignment but degrades subject consistency.:
4.4.2. Hyperparameter Analysis
We conduct additional ablation experiments to analyze the impact of similarity threshold $\tau$ and cycle consistency threshold $\delta$ on semantic matching. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack, Fig. 9).
We further verify the effectiveness of dropout in mitigating structural overfitting. As shown in Fig. 10, moderate dropout enhances pose flexibility while preserving subject identity, whereas excessive dropout intensities ($\omega>0.6$) erode identity retention. These findings highlight the necessity of balancing pose flexibility and detail preservation through parameter tuning.
Figure 11. Comparison between DreamBooth and our method on FLUX.1. Without computation-intensive fine-tuning, our method achieves performance comparable to DreamBooth and even surpasses it in capturing visual details such as text and patterns.
Figure 12. Generation results of (a) Stable Diffusion (b)Stable Diffusion with DreamMatcher (c) DreamBooth on Stable Diffusion (d) DreamBooth on Stable Diffusion with DreamMatcher (e) Ours.
4.5. Comparison with DreamBooth
We compare FreeGraftor with DreamBooth, a widely-used tuning-based method. Due to limitations in computational resources, we implemented DreamBooth using LoRA instead of full parameter fine-tuning on FLUX.1-Dev, with a LoRA rank of 16.
As shown in Fig. 11, although fine-tuned for several tens of minutes on the reference images of each subject, DreamBooth still struggles to fully reproduce the visual details of the reference subject (e.g., text and patterns). Additionally, fine-tuning on limited samples can lead to overfitting of the background, compromising its ability to follow textual semantics. In contrast, without resource-intensive fine-tuning, our method achieves pixel-level detail preservation and faithful text alignment.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
DreamBooth 0.9261 0.8776 0.3123 1.3232
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 3. Quantitative Comparison of DreamBooth and our method on FLUX.1-Dev.
Furthermore, we quantitatively compared FreeGraftor with DreamBooth using the same experimental setup (Sec. 4.1.3). Table 3 shows our approach delivers significant performance advantages beyond just efficiency gains. FreeGraftor resolves the fidelity-efficiency dilemma in subject-driven generation, going beyond merely alleviating the limitations of training-free methods.
4.6. Comparison with DreamMatcher
The most comparable approach to ours is likely DreamMatcher(Nam et al., 2024), yet the two remain fundamentally distinct. Firstly, DreamMatcher is not a standalone subject-driven generation method but rather a plugin. It requires integration with tuning-based approaches (e.g., DreamBooth), which involve time-consuming and resource-intensive test-time optimization. In contrast, our method is a standalone subject-driven generation pipeline without any fine-tuning or training. Secondly, DreamMatcher adopts a direct attention feature replacement mechanism for identity injection, which tends to introduce artifacts along subject boundaries. In contrast, our approach leverages cross-image attention with shared spatial information, employing a more refined methodology that effectively mitigates image distortion.
Since DreamMatcher is tailored specifically for Stable Diffusion 1.5 and cannot be directly migrated to other network architectures, we implemented DreamMatcher on Stable Diffusion 1.5 following the official code. As show in Fig. 12, when employed in conjunction with DreamBooth, DreamMatcher effectively rectifies erroneous attribute binding(e.g., the yellow number ”3” on the clock), thereby enhancing the fidelity of reconstructed visual details. However, when applied directly to Stable Diffusion, DreamMatcher is unable to fully reconstruct the subject from the reference image, owing to a lack of prior structural knowledge of the target subject. Additionally, some images generated by DreamMatcher exhibited noticeable distortions (e.g., green and blue artifacts around the backpack).
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
Stable Diffusion + DreamMatcher 0.7830 0.5336 0.3232 0.6387
DreamBooth + DreamMatcher 0.8912 0.7918 0.3021 0.1767
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 4. Quantitative Comparison of DreamMatcher and our method.
Quantitative comparisons under identical settings (Sec. 4.1.3) confirm our superiority, as shown in Table 4. Our method achieves significant performance gains through two key designs. First, the structure-consistent noise initialization produces reasonable image layouts while faithfully preserving reference geometry, which ensures robust semantic matching and detail retention. Second, our position-constrained attention fusion mechanism enables harmonious integration of features, resulting in more natural-looking outputs. Ablation studies (Sec. 4.4.1) confirm the effectiveness of these core components.
5. Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly transfers the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control while requiring no training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
Acknowledgements.
This work was supported by Beijing Natural Science Foundation (L247010), NSFC (No.61906018).